# Quantifying Memorization and Retriever Performance in Retrieval-Augmented Vision-Language Models

### v1

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable capabilities in question answering (QA), but metrics for assessing their reliance on memorization versus retrieval remain underdeveloped. Moreover, while finetuned models are state-of-the-art on closed-domain tasks, general-purpose models like GPT-4o exhibit strong zero-shot performance. This raises questions about the trade-offs between memorization, generalization, and retrieval. In this work, we analyze the extent to which multimodal retrieval-augmented VLMs memorize training data compared to baseline VLMs. Using the WebQA benchmark, we contrast finetuned models with baseline VLMs on multihop retrieval and question answering, examining the impact of finetuning on data memorization. To quantify memorization in end-to-end retrieval and QA systems, we propose several proxy metrics by investigating instances where QA succeeds despite retrieval failing. Our results reveal the extent to which finetuned models rely on memorization. In contrast, retrieval-augmented VLMs have lower memorization scores, at the cost of accuracy (72% vs 52% on WebQA test set). As such, our measures pose a challenge for future work to reconcile memorization and generalization in both Open-Domain QA and joint Retrieval-QA tasks.
Quantifying Memorization and Retriever Performance in Retrieval-Augmented Vision-Language Models
@@ 1 Introduction
The increasing reliance on LLMs for multimodal tasks across far-reaching sectors such as healthcare, finance, and manufacturing underscores the need to assess the accuracy and reliability of the information they generate. Vision-Language Models (VLM) have achieved state-of-the-art (SoTA) performance on Visual Question-Answering (VQA) benchmarks, and these models often utilize Retrieval-Augmented Generation (RAG) to maintain factual accuracy and relevance in a dynamic information environment. However, this has led to uncertainty in the information the LLM bases its answer on, as it may choose between parametric memory and retrieved sources. When models rely on memorized information instead of dynamically retrieving information, they may inadvertently propagate outdated or incorrect information, causing serious legal and ethical risks and undermining trust and reliability in AI systems (Huang et al., 2023).
@@
Despite these concerns, the way that Vision-Language models (VLMs) memorize and retrieve information, particularly in complex multimodal tasks, remains under-explored. Current research often focuses on either the general capabilities of large language models (LLMs) or the specialized retrieval mechanisms in retrieval augmented generation systems (RAG) (Ram et al., 2023; Chen et al., 2022b; Liu et al., 2023). Particularly in the context of multimodal retrieval and multihop reasoning, few studies analyze the tradeoff between finetuning for specialized tasks and zero-shot prompting for general-purpose vision-language capabilities. A lack of consensus on how to approach this tradeoff motivates the development of measures to quantify reliance on parametric memory, as well as metrics for quantifying the potential performance impact of extending LLMs with RAG systems.
@@
To address this gap, we investigate how multimodal QA models balance accuracy with memorization on the WebQA benchmark. We compare finetuned multimodal systems against zero-shot VLMs, analyzing how retrieval performance influences QA accuracy. In particular, we focus on cases where retrieval fails, allowing us to measure reliance on parametric memory through two proposed metrics—the Parametric Proxy Rate (PPR) which quantifies how much model accuracy is influenced by retrieval quality, contrasting performance in best-case versus worst-case retrieval scenarios, and the Unsupported Correctness Rate (UCR) which measures how often correct QA responses are generated when the retriever fails, providing a proxy for memorization.
@@
To enable this analysis, we make several methodological contributions. For the finetuned QA models, we investigate Vision-Transformer (ViT) architectures, which allow for multihop reasoning over multiple sources. To investigate the impact of retrieval performance on trained LMs, we propose a variable-input Fusion-in-Decoder (FiD) model Tanaka et al. (2023); Suhr et al. (2018), building upon the VoLTA architecture (Pramanick et al., 2022). For the zero-shot case, we build upon previous research on In-Context Retrieval (Ram et al., 2023) by demonstrating that LLMs such as GPT-4o are capable of performing the final ranking step of the retrieval process. In doing so, we find that GPT-4o, a general-purpose LLM, achieves SoTA performance on the WebQA task, outperforming existing finetuned RAG models by a significant margin (7% higher accuracy).
Crucially, our results reveal that while retrieval-augmented models reduce memorization, the training paradigm plays an important role. Finetuned models exhibit higher reliance on parametric memory, whereas zero-shot RAG approaches have lower memorization scores at the cost of accuracy. This suggests that while retrieval modules may mitigate the risks associated with outdated or incorrect information, SoTA performance requires that they be coupled with specialized QA models. Our memorization measures contribute to the development of transparent and reliable AI systems, particularly in applications where the sourcing of up-to-date, factual information is critical.
@@ 2.3 The Parametric Effect
There is a wealth of research on reliance on parametric memory for unimodal retrieval and QA tasks (Galway et al., 2024; Xu et al., 2024; Longpre et al., 2022; Neeman et al., 2022; Hong et al., 2024; Chen et al., 2022a). Here, the entity replacement framework (Longpre et al., 2022; Neeman et al., 2022) is used to invalidate parametric memory by explicitly crafting knowledge conflicts between input sources and parametric memory (Xu et al., 2024; Hong et al., 2024; Chen et al., 2022a). As such, these studies guarantee that manipulated input sources no longer entail the expected labels. In contrast, we do not make the same guarantees, and our proxy measures are premised upon the key assumption that incorrectly retrieved sources do not entail the correct answer. Our focus is on developing proxy metrics for the parametric effect that do not require such involved source manipulation processes, so extending the entity replacement framework to the multimodal setting is left as the subject of future work.
3 Datasets
WebQA
The WebQA dataset Chang et al. (2022) is designed with a two-step approach in mind; retrieval followed by QA. First, given the question Q and all sources $S$, we retrieve the set of relevant sources, $S^{\prime}$. Using these sources we then generate an answer $A^{\prime}$. The following is passed to the QA classifier:
@@ 3 Datasets
We include only those questions that either one (n = 12,027) or two (n = 9,438) image sources. For a breakdown of question categories and their keywords, see subsection A.6 in the appendix.
As opposed to WebQA, open-domain VQA tasks such as OK-VQA (Marino et al., 2019) and HotpotQA (Yang et al., 2018) do not provide candidate sources $S$ and source labels $S^{*}$. As such, the memorization metrics proposed are incompatible with these tasks (see section 5). Moreover, while we do evaluate QA model performance on NLVR2 (Suhr et al., 2018) and VQAv2 (Goyal et al., 2017), these tasks do not have a retrieval step and so these results are instead presented in the appendix (see subsection A.1).
@@ 4 Methodology
As WebQA is a joint retrieval and QA task, we develop several QA methods and retrieval methods separately. For the final end-to-end task, we evaluate the combination of the each retrieval method with the best-performing QA model, as determined during an initial QA model selection phase. This enables us to investigate the factors involved in both the model design and the model finetuning process that may influence how the parametric effect manifests itself.
@@ 4.1 Question Answering
We select the VoLTA framework as the skeleton for encoding joint text and image representations Pramanick et al. (2022). VoLTA uses Swin-Base Liu et al. (2021) and RoBERTa-Base Liu et al. (2019) as respective visual and textual encoders and we adopt the same encoder choices. Swin Transformer introduced a hierarchical Transformer design with a shifted windows scheme, which computes the self-attention layer over non-overlapping local windows and allows connection by shifting between windows. We jointly encode each image source returned by the retriever with the query and concatenate the resulting embeddings together before sending them to the MLP classifier to predict the keyword answer label. To handle variable input sequences during classification, we pad single image sources with blank images so that all inputs sent into the classifiers have two images. We call this model MultiHop-VoLTA (MH-VoLTA).
@@ 5.1 Unsupported Correctness Rate
UCR(R,M) $=\frac{\text{Acc}_{M}==1\cap\text{Recall}_{R}=0}{\text{Recall}_{R}=0}$ (7)
@@ 5.2 Oracle-Normalized Retrieval Scores
These metrics are Retriever Potential Attainment (RPA, higher is better) and Parametric Proxy Rate (PPR, lower is better).
@@
RPA quantifies the potential that a retriever has realized when used in a given end-to-end retrieval QA system. The upper bound (1) is given by same QA system’s accuracy with oracle sources ($\text{Acc}_{M}(\text{oracle})$). The lower bound (0) is given by the random negative source retriever ($\text{Acc}_{M}(\text{random})$), which always retrieves incorrect sources. We apply random-oracle scaling, where $Acc_{M}(R)$ denotes the accuracy of QA model M, given sources from retriever R:
@@
Question Complexity Analysis Metrics
To identify correlations between the complexity of the question, retrieval recall, and QA accuracy, we apply three separate measures to the input questions; Word Count, Flesch-Kincaid Grade Level (Flesch, 2007), and Gunning-Fog Index (Gunning, 1952)(subsection A.5).
@@ 6 Results
First, we experiment with multiple QA models to determine which generalized LLMs and specialized, finetuned models should be selected for the joint retrieval and QA task. Then, we experiment with different combinations of retrieval systems and chosen QA models for the joint task. Finally, we analyze how performance on the retrieval task impacts QA accuracy and investigate the relationship between question complexity, performance, and model memorization.
@@ 6.1 Model Selection
(b) Across experiments, lower retrieval recall impacts QA accuracy as distractor sources hurt performance.
@@ 6.2 Impact of Retrieval on QA
To understand how retrieval and QA systems interact, we investigate the reliance of the QA task on retrieval correctness. We find that when GPT-4o correctly retrieves the relevant sources (retrieval recall is high), it has a 59% QA accuracy rate. Conversely, if GPT-4o does not retrieve the correct sources, the accuracy rate is reduced to 26% (1(a)). We also find that QA performance drops as the number of retrieved distractors increases and retrieval recall falls, showing that poor retrieval performance adversely affects QA (1(b)). This is to say that the QA task is heavily dependent on retrieval performance. However, there do exist correctly answered questions for which incorrect sources are retrieved, and these samples form the basis for the UCR measure of the parametric effect (1(a)).
@@ 6.4 Finetuning and UCR
However the effect of retrieval on UCR is not negligible, and we find that for a given QA model, UCR increases as retrieval recall increases; i.e. for each model M $\text{UCR}(\text{Rand},M)<\text{UCR}(\text{Clip},M)<\text{UCR}(\text{ClipDPR},M)$ (Table 2). This implies that as the retriever improves, the QA model is more successful on samples that retrieval fails on. This paradox is explained by inaccuracies in the source labels—‘distractor’ sources often provide enough evidence for the QA model to answer correctly. Rather than exposing memorization, this reveals an underlying issue with the source labels in the WebQA dataset, and as such, these measures can be adapted to evaluate the correctness of the joint retrieval-QA benchmarks.
Table 2: UCR ($P(Q_{1}|R_{0})$) for each retriever R and QA model pairing, alongside retrieval recall. CDPR denotes Clip-DPR, Q-FT is Qwen2-FT.
R recall Qwen Q-FT MHV GPT4
@@
6.5 Question Complexity
We observe and report interesting relationships between query complexity and retrieval and QA performance. We find that the accuracy of the in-context GPT-4o retriever is related to question complexity (Figure 3). The more complex the question in terms of word count, Flesch-Kincaid Grade, or Gunning Fog Index, the lower the QA performance (Figure 3). Conversely, increasing query complexity improves GPT-4o’s retrieval ability, where the additional complexity provides information on source relevancy  [^1] We note that this opposing relationship between retrieval and QA performance is contrary to the finding that the QA task is heavily dependent on retrieval performance (1(b)). The impact of query complexity on task performance is strong enough to overcome this general principle.. However, this relationship does not hold for the finetuned UniVL-DR retriever, where question complexity has little effect on retrieval recall or QA accuracy (Figure 4). As such, systems that rely on "in-context" retrieval using GPT-4o are limited by query complexity, but approaches that utilize finetuned retrievers are not.
@@ 7 Discussion
While QA performance is generally predicated upon retrieval success (1(b)), there are many cases where retrieval fails and QA succeeds (1(a)). These cases form the basis of our quantitative metrics, the Unsupported Correctness Rate (UCR; see Table 2) and Parametric Proxy Rate (PPR; see Figure 2), and with these measures we show that external retrievers significantly reduce the reliance of VLMs on parametric memory. This reduction in memorization not only preserves model flexibility but also mitigates the over-specialization common in finetuned systems. However, despite GPT-4o obtaining state-of-the-art performance on the WebQA benchmark using this approach (Table 1)222“Anon_ACL_ARR_Feb25” @ WebQA leaderboard, for less powerful VLMs such as Qwen2 the decrease in memorization associated with not finetuning the model (PPR: 0.77 -> 0.53) comes at the cost of model accuracy (QA accuracy: 70% -> 52%).
Building on previous research in retrieval-augmented generation and multimodal question answering, our study explicitly quantifies the balance between memorization and retrieval. Our findings reveal that in-context retrieval can effectively substitute for finetuning in multimodal QA tasks. However, this approach is currently limited by the complexity of the question and task (Figure 3). The strong performance of general-purpose VLMs when augmented with finetuned retrievers suggests that general-purpose VLMs are capable of reasoning over retrieved documents and images, without being affected by question complexity (Figure 4).
This analysis also reveals a paradoxical relationship between retriever recall and UCR that highlights a potential annotation issue within the dataset, prompting a reevaluation of how retrieval-QA benchmarks are constructed (Table 2). As such, our approach is subject to certain limitations. Annotation inconsistencies may affect the reliability of UCR as a sole indicator of retrieval quality. Moreover, measures that more directly evaluate model attention are feasible, and may incorporate visual attention models such as Grad-CAM (Selvaraju et al., 2020) or object detection methods (Ravi et al., 2024). This line of work can take inspiration from a wealth of research on provoking parametric responses from unimodal LLMs, where the entity replacement methods (Longpre et al., 2022; Neeman et al., 2022) create knowledge conflicts between input sources and parametric memory (Xu et al., 2024; Hong et al., 2024; Chen et al., 2022a). By explicitly crafting knowledge conflicts, extending these methods to the multimodal setting will allow memorization analysis to move beyond the assumption that incorrectly retrieved sources should result in incorrect answers from the VQA model.
@@
Despite this assumption, our measures reveal an interesting interplay between retrieval and memorization. By providing insights into end-to-end retrieval and QA systems, UCR can highlight when models are over-reliant on parametric memory. High Retriever Attainment scores (RPA) across Qwen2 and GPT-4o experiments demonstrate that general-purpose VLMs can utilize finetuned retrievers (Figure 2), drawing the need for domain-specific finetuning into question. This work points towards In-Context Retrieval as a particularly promising direction for future research in multimodal systems, if the limitation of question complexity can be addressed.
Our findings suggest concrete directions for improving retrieval-augmented QA systems. First, incorporating Unsupported Correctness Rate into the evaluation of retrieval tasks can highlight annotation issues, as in WebQA (Table 2). Second, integrating retrieval-aware finetuning strategies—–where LLMs learn when to prioritize retrieved content rather than rely on parametric knowledge (Labruna et al., 2024)—–may help further reduce memorization while maintaining accuracy. Third, extending our memorization analysis to open-domain multimodal tasks would help validate our findings while providing new insights into retrieval dynamics in unrestricted, real-world settings. Addressing these challenges will contribute to the development of AI systems that are not only more generalizable but also more interpretable, reducing the risks of memorization-driven hallucination.
@@
Overall, our methodology provides a framework for measuring and mitigating memorization in retrieval-augmented systems, offering new ways to evaluate the quality of retrieval-QA datasets. By quantifying reliance on parametric memory, researchers can better assess the trade-offs within finetuning and retrieval, thereby guiding the development of models that balance generalization with task-specific performance. As retrieval-augmented models continue to scale, our findings highlight the need for robust retrieval evaluation and memorization control to ensure safe, effective, and adaptable AI systems. This work lays a foundation for future research into retrieval-aware training strategies, multimodal in-context learning, and more interpretable AI architectures that can integrate external knowledge while minimizing unintended memorization.

### v2

# Quantifying Memorization and Parametric Response Rates in Retrieval-Augmented Vision-Language Models

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable capabilities in question answering (QA), but metrics for assessing their reliance on memorization versus retrieval remain underdeveloped. Moreover, while finetuned models are state-of-the-art on closed-domain tasks, general-purpose models like GPT-4o exhibit strong zero-shot performance. This raises questions about the trade-offs between memorization, generalization, and retrieval. In this work, we analyze the extent to which multimodal retrieval-augmented VLMs memorize training data compared to baseline VLMs. Using the WebQA benchmark, we contrast finetuned models with baseline VLMs on multihop retrieval and question answering, examining the impact of finetuning on data memorization. To quantify memorization in end-to-end retrieval and QA systems, we propose several proxy metrics by investigating instances where QA succeeds despite retrieval failing. In line with existing work, we find that finetuned models rely more heavily on memorization than retrieval-augmented VLMs, and achieve higher accuracy as a result (72% vs 52% on WebQA test set). Finally, we present the first empirical comparison of the parametric effect between text and visual modalities. Here, we find that image-based questions have parametric response rates that are consistently 15-25% higher than for text-based questions in the WebQA dataset. As such, our measures pose a challenge for future work, both to account for differences in model memorization across different modalities and more generally to reconcile memorization and generalization in joint Retrieval-QA tasks.
Quantifying Memorization and Parametric Response Rates in Retrieval-Augmented Vision-Language Models
@@ 1 Introduction
The increasing reliance on LLMs for multimodal tasks across far-reaching sectors such as healthcare, finance, and manufacturing underscores the need to assess the accuracy and reliability of the information they generate. Vision-Language Models (VLM) have achieved state-of-the-art (SoTA) performance on Visual Question-Answering (VQA) benchmarks, and these models often utilize Retrieval-Augmented Generation (RAG) to maintain factual accuracy and relevance in a dynamic information environment. However, this has led to uncertainty in the information the LLM bases its answer on in situations where it may choose between parametric memory and retrieved sources. When models rely on memorized information instead of dynamically retrieving information, they may inadvertently propagate outdated or incorrect information, causing serious legal and ethical risks and undermining trust and reliability in AI systems (Huang et al., 2023).
@@
Despite these concerns, the way that Vision-Language models (VLMs) memorize and retrieve information, particularly in complex multimodal tasks, remains under-explored. Instead, survey studies on parametric knowledge conflicts have found that existing research is focused on reasoning capabilities of unimodal large language models (LLMs) and retrieval augmented generation systems (RAG) (Xu et al., 2024a). Particularly in the context of multimodal retrieval and multihop reasoning, few studies analyze the tradeoff between finetuning for specialized tasks and zero-shot prompting for general-purpose vision-language capabilities. A lack of consensus on how to approach this tradeoff motivates the development of measures to quantify reliance on parametric memory, as well as metrics for quantifying the potential performance impact of extending LLMs with RAG systems.
@@
To address this gap, we investigate how multimodal QA models balance accuracy with memorization on the WebQA benchmark. We compare finetuned multimodal systems against zero-shot VLMs, analyzing how retrieval performance influences QA accuracy. In particular, we focus on cases where retrieval fails, allowing us to measure reliance on parametric memory through two proposed metrics—the Parametric Proxy Rate (PPR) which quantifies how much model accuracy is influenced by retrieval quality, contrasting performance in best-case versus worst-case retrieval scenarios, and the Unsupported Correctness Rate (UCR) which measures how often correct QA responses are generated when the retriever fails, providing a proxy for memorization. Figure 1 gives an overview of how these measures are derived for a joint retrieval-QA task.
Figure 1: PPR and UCR metrics are derived from the interaction between retrieved sources and QA model.
@@
To enable this analysis, we make several methodological contributions. For the finetuned QA models, we investigate Vision-Transformer (ViT) architectures, which allow for multihop reasoning over multiple sources. To investigate the impact of retrieval performance on trained LMs, we propose a variable-input Fusion-in-Decoder (FiD) model Tanaka et al. (2023); Suhr et al. (2018), building upon the VoLTA architecture (Pramanick et al., 2022). For the zero-shot case, we build upon previous research on In-Context Retrieval (Ram et al., 2023) by demonstrating that LLMs such as GPT-4o are capable of performing the final ranking step of the retrieval process. In doing so, we find that GPT-4o, a general-purpose LLM, achieves SoTA performance on the WebQA task, outperforming existing finetuned RAG models by a significant margin (7% higher accuracy). Crucially, our results reveal that while retrieval-augmented models reduce memorization, the training paradigm plays an important role. Finetuned models exhibit higher reliance on parametric memory, whereas zero-shot RAG approaches have lower memorization scores at the cost of accuracy. While retrievers the improve performance zero-shot VLMs to some degree, as is the case for unimodal systems there is yet no silver bullet in the tradeoff between model generalization and specialization.
Finally, we investigate differences in parametric response rates between text-based and image-based questions from the WebQA dataset. We find that models are capable of answering image-based questions based on parametric knowledge 15-25% more often than they are for text-based sources. In many cases, this means that parametric responses are twice as likely for image-based questions as for text-based ones. Moreover, this result is consistent regardless of the retriever used or whether the QA model was finetuned or not. This finding represents the key empirical contribution from this work—not only is this the first work to measure the parametric effect over image sources, but to the best of our knowledge, it is also the first to present empirical results comparing model memorization tendencies across different modalities. Our findings suggest that the parametric effect may be more pronounced for visual tasks. Future work to validate these findings on additional datasets and problem domains is warranted. We hope that our analysis of model memorization motivates the development of transparent and reliable multimodal AI systems, particularly in applications where the sourcing of up-to-date, factual information from multimodal sources is critical.
@@ 2.3 The Parametric Effect
There is a wealth of research on reliance on parametric memory for unimodal QA tasks (Galway et al., 2024; Xu et al., 2024b; Longpre et al., 2022; Neeman et al., 2022; Hong et al., 2024; Chen et al., 2022a). Here, the entity replacement framework (Longpre et al., 2022; Neeman et al., 2022) is used to invalidate parametric memory by explicitly crafting knowledge conflicts between input sources and parametric memory (Xu et al., 2024b; Hong et al., 2024; Chen et al., 2022a). As such, these studies guarantee that manipulated input sources no longer entail the expected labels, and focus on evaluating LLMs in isolation without using retrieval systems.
In contrast, we do not make the same guarantees, and our proxy measures are premised upon the key assumption that incorrectly retrieved sources do not entail the correct answer. Our focus is on developing proxy metrics for the parametric effect that do not require such involved source manipulation processes. Rather, building upon prior work on unimodal LLMs (Soudani et al., 2024), these metrics compare the performance of finetuned VQA models with RAG systems.
3 WebQA Dataset
The WebQA dataset Chang et al. (2022) uses a two-step design; retrieval followed by QA. First, given the question Q and all sources $S$, we retrieve the set of relevant sources, $S^{\prime}$. Using these sources we then generate an answer $A^{\prime}$. The following is passed to the QA classifier:
@@ 3 WebQA Dataset
We include only those questions that require either one (n = 12,027) or two (n = 9,438) image sources. For a breakdown of question categories and their keywords, see 8 in the appendix. The remaining questions use only text sources (n = 20,267). Our final analysis of unimodal vs multimodal parametric effects uses this portion of the dataset to evaluate memorization on text sources.
As opposed to WebQA, open-domain VQA tasks such as OK-VQA (Marino et al., 2019) and HotpotQA (Yang et al., 2018) do not provide candidate sources $S$ and source labels $S^{*}$, and as a result are incompatible with or measures (see section 5). Moreover, while we do evaluate model performance on VQA datasets (NLVR2 (Suhr et al., 2018) and VQAv2 (Goyal et al., 2017)), these tasks lack a retrieval step and so are only useful for QA model selection (see subsection A.1).
@@ 4 Methodology
As WebQA is a joint retrieval and QA task, we develop several QA methods and retrieval methods separately. Using the best-performing QA model, we then evaluate end-to-end retrieval and VQA performance and investigate the factors that affect the parametric effect.
@@ 4.1 Question Answering
We select the VoLTA framework as the skeleton for encoding joint text and image representations Pramanick et al. (2022). VoLTA uses Swin-Base Liu et al. (2021) and RoBERTa-Base Liu et al. (2019) as respective visual and textual encoders and we adopt the same encoder choices. We jointly encode each image source returned by the retriever with the query and concatenate the resulting embeddings together before sending them to the MLP classifier to predict the keyword answer label. To handle variable input sequences during classification, we pad single image sources with blank images so that all inputs sent into the classifiers have two images. We call this model MultiHop-VoLTA (MH-VoLTA).
@@ 5.1 Unsupported Correctness Rate
UCR(R,M) $=\frac{\text{Acc}_{M}==1\cap\text{Recall}_{R}==0}{\text{Recall}_{R}==0}$ (7)
@@ 5.2 Oracle-Normalized Retrieval Scores
RPA quantifies the potential that a retriever has realized when used in a given end-to-end retrieval QA system. The upper bound (1) is given by same QA system’s accuracy with oracle sources ($\text{Acc}_{M}(\text{oracle})$). The lower bound (0) is given by the random negative source retriever ($\text{Acc}_{M}(\text{random})$), which always retrieves incorrect sources. Note that $\text{Acc}_{M}(\text{random})$ is similar to the Free Success Rate metric (Lin and Byrne, 2022), which represents the QA model’s accuracy with no retrieved sources. We apply random-oracle scaling, where $Acc_{M}(R)$ denotes the accuracy of QA model M, given sources from retriever R:
@@ 6 Results
First, we experiment with multiple QA models to determine which generalized LLMs and specialized, finetuned models should be selected for the joint retrieval and QA task. Then, we analyze how performance on the retrieval task impacts QA accuracy and experiment with different combinations of retrieval systems and chosen QA models for the joint task. Finally, we investigate the effect of finetuning on model memorization and compare parametric response rates over WebQA image-based and text-based questions.
@@ 6.1 Model Selection
(b) Retrieving distractor sources decreases QA accuracy.
Figure 2: Across experiments, recall impacts QA.
@@ 6.2 Impact of Retrieval on QA
To understand how retrieval and QA systems interact, we investigate the reliance of the QA task on retrieval correctness (Figure 2). We find that when GPT-4o correctly retrieves the relevant sources through In-Context Retrieval, it has a 59% QA accuracy rate. If GPT-4o In-Context Retrieval fails to retrieve the correct sources, the accuracy rate is reduced to 26% (2(a)). We also find that QA performance drops as the number of retrieved distractors increases and retrieval recall falls, showing that poor retrieval performance adversely affects QA (2(b)). This is to say that the QA task is heavily dependent on retrieval performance. However, there do exist correctly answered questions for which incorrect sources are retrieved, and these samples form the basis for the UCR measure of the parametric effect (2(a)).
In addition, we contribute error analysis in the appendix that show the differences between In-Context Retrieval using GPT-4o and dense retrieval methods such as UniVL-DR (subsection A.5). In particular, we find that systems that rely on "in-context" retrieval using GPT-4o are limited by query complexity, but approaches that utilize dense retrievers are not. As query complexity increases, In-Context Retrieval degrades QA accuracy (Figure 6), but dense retrievers do not (Figure 7).
@@ 6.4 Finetuning and UCR
However, the effect of retrieval on UCR is not negligible, and we find that for a given QA model, UCR increases as retrieval recall increases; i.e. for each model M $\text{UCR}(\text{Rand},M)<\text{UCR}(\text{Clip},M)<\text{UCR}(\text{ClipDPR},M)$ (Table 2). This implies that as the retriever improves, the QA model is more successful on samples that retrieval fails on. This paradox is explained by inaccuracies in the source labels—distractor sources often provide enough context for the QA model to answer correctly. Rather than exposing memorization, this reveals an underlying issue with the source labels in the WebQA dataset, and as such, these measures can be adapted to evaluate the correctness of the joint retrieval-QA benchmarks.
Retri- recall Unsupported Correctness Rate
ver R Qwen Q-FT MHV GPT4
@@
Table 2: UCR ($P(Q_{1}|R_{0})$) for each retriever R and QA model M, alongside retrieval recall. CDPR denotes Clip-DPR, Q-FT is Qwen2-FT.
6.5 Multimodal vs Unimodal Memorization
Finally, using the developed metrics and the finetuned and non-finetuned VLM models, we investigate the differences between textual and visual parametric knowledge. Figure 4 reveals that parametric responses are more prevalent for webqa image questions than for webqa text questions. While finetuning results in a minor increase of parametric response rates for the text modality, these rates increase dramatically for the visual modality after finetuning. For example, while UCR on text-based questions increases by $\sim 10\%$ with finetuning, it increases by $\sim 20\%$ for image-based questions. In addition, parametric responses are as much as twice as likely when models are presented with image sources, and these results are consistent across the metrics and models used. For example, for Qwen2-FT, $\text{PPR}_{img}=0.77$ while $\text{PPR}_{txt}=0.4$, and when Qwen is paired with the UniVL-DR retriever, the end-to-end system has $\text{UCR}_{img}=0.44$ and $\text{UCR}_{txt}=0.22$. These results highlight a modality-based disparity in how memorization manifests in QA models.
(a) $\text{PPR}(\text{image})>\text{PPR}(\text{text})$
(b) $\text{UCR}(\text{image})>\text{UCR}(\text{text})$
Figure 4: Higher implies greater levels of memorization. Across all metrics, QA models, and retrievers tested, parametric responses are more prevalent for webqa image questions than for webqa text questions.
@@ 7 Discussion
While QA performance is generally predicated upon retrieval success (2(b)), there are many cases where retrieval fails and QA succeeds (2(a)). These cases form the basis of our quantitative metrics, the Unsupported Correctness Rate (UCR; see Table 2) and Parametric Proxy Rate (PPR; see Figure 3), and with these measures we show that external retrievers significantly reduce the reliance of VLMs on parametric memory. This not only preserves model flexibility but also mitigates the over-specialization common in finetuned systems. However, despite GPT-4o obtaining state-of-the-art performance on the WebQA benchmark using this approach (Table 1)111“Anon_Feb25” @ WebQA leaderboard, for less powerful VLMs such as Qwen2 the decrease in Parametric Proxy Rate associated with not finetuning the model (PPR: 0.77 -> 0.53) comes at the cost of model accuracy (QA accuracy: 70% -> 52%).
In interpreting these results, it is important to note that UCR and PPR are proxy measures based upon the key assumption that incorrectly retrieved sources should result in incorrect answers from the VQA model. While we can guarantee that distractor sources from the random retriever provide no useful information to the model (i.e. they are randomly sampled negatives), fully addressing this assumption requires modifying the image sources to invalidate the original answer or label. Along these lines, a recent line of research uses constrained image generation to create knowledge conflicts between image sources and parametric memory (Carragher et al., 2025). This builds on research into provoking parametric responses from unimodal LLMs, where the entity replacement methods (Longpre et al., 2022; Neeman et al., 2022) create knowledge conflicts between text sources and parametric memory (Xu et al., 2024b; Hong et al., 2024; Chen et al., 2022a). Entity replacement frameworks for VLMs (Carragher et al., 2025) can make use of object detection (Ravi et al., 2024) and visual attention models (Selvaraju et al., 2020) to allow parametric analysis to move beyond the incorrectness assumption.
@@
Despite this assumption, our measures reveal an interesting interplay between retrieval and parametric responses. By providing insights into end-to-end retrieval and QA systems, UCR can highlight when models are over-reliant on parametric memory. High Retriever Attainment scores (RPA) across Qwen2 and GPT-4o experiments demonstrate that general-purpose VLMs can utilize finetuned retrievers (Figure 3), drawing the need for domain-specific finetuning into question. This work points towards In-Context Retrieval as a particularly promising direction for future research in multimodal systems, if the limitation regarding question complexity can be addressed (A.5).
Crucially, while image manipulation methods are not subject to the incorrectness assumption that the proxy metrics proposed here are, they are not applicable to text sources. Our proxy measures allow for the comparison of how the parametric effect manifests across different modalities in multimodal language models. Herein, we find that parametric responses may be more prominent over image sources as opposed to text sources (Figure 4). Given that the parametric effect is, as of yet, understudied in multimodal setting (Xu et al., 2024a), to the best of the authors knowledge this is the first comparison of parametric effects between modalities. Our finding motivates incorporating the wealth of parametric research on unimodal models into the multimodal domain (Carragher et al., 2025).
@@
Overall, our methodology provides a framework for measuring and mitigating memorization in retrieval-augmented systems, offering new ways to evaluate the quality of retrieval-QA datasets. Future work can apply this framework to improve retrieval-aware finetuning strategies, where LLMs learn when to prioritize retrieved content rather than rely on parametric knowledge (Labruna et al., 2024). Extending our parametric analysis to open-domain multimodal tasks would provide new insights into retrieval dynamics in unrestricted, real-world settings. Finally, by quantifying reliance on parametric memory, researchers can better assess the trade-offs within finetuning and retrieval per modality, thereby guiding the development of multimodal models that balance generalization with task-specific performance. As retrieval-augmented VLMs continue to scale, our findings highlight the need for multimodal evaluation of parametric responses to ensure safe, effective, and adaptable AI systems.
7.1 Limitations
The UCR analysis presented in 4(b) is subject to limitations based on the design of the metric. Specifically, if certain question categories have higher retrieval recall, the number of incorrectly retrieved samples for that category will be low. A study of UCR across question categories, including suitable confidence intervals, is warranted.
Our analysis also reveals a paradoxical relationship between retriever recall and UCR that highlights a potential annotation issue within the dataset, prompting a reevaluation of how retrieval-QA benchmarks are constructed (Table 2). While this means that UCR can be used in the evaluation of retrieval tasks to highlight potential false negative annotation issues, as in WebQA (Table 2), annotation inconsistencies in turn affect the reliability of UCR as a sole indicator of retrieval quality. As such, our findings should be validated on additional VQA datasets. To facilitate this, widely used VQA datasets could be augmented with retrieval tasks, such that joint retrieval-QA systems may be evaluated on them.
Finally, given the rapid pace of research in the multimodal space, the WebQA dataset may have already been incorporated into the training data of the VLMs investigated here. While WebQA is not specifically listed among the Qwen2-VL training materials (Wang et al., 2024), many similar datasets are. For models that have been pretrained on the same joint retrieval-QA dataset used to operationalize the measures proposed here, the measures may be more indicative of verbatim memorization, where model output exactly matches the dataset labels. In contrast, our analysis is targeted at the parametric effect, which is a preference for models to reason over parametric knowledge instead of input sources. Disentangling verbatim memorization from the parametric effect is a good avenue for future research, as it represents a more dramatic failure case of model generalization.